Find the volume of a sphere with a radius measuring 4.5 inches.
I have tried to solve multiple times and my book isn't explaining it well, need another source.

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

Find the volume of a sphere with a radius measuring 4.5 inches.
---
V = (4/3)(pi)r^3
---
V = (4/3)(pi)1.5^3
---
V = 4.5(pi) in^3
---
V is approximately 14.14 in^3
====================================
Cheers,
Stan H.
===============

Answer by solver91311(24713)



The volume of a sphere is four thirds times times the radius cubed.

Multiply 4.5 times 4.5. Multiply the result times 4.5. That gives you the radius cubed.

Multiply the radius cubed by 4 and divide the result by 3. That gives you the radius cubed times four thirds.

Then you need to decide whether you want the exact answer or a numeric approximation.

If you want the exact answer, take the result of your calculation, and write that next to and you are done. If you want a numeric approximation, then chose an approximation of suitable for the precision you desire in your answer and then multiply that times the result of the calculation above, rounding the result appropriately. Either way, remember to express your answer in terms of cubic inches because your radius measurement was given in inches and you had to cube that number in the process of calculating your answer.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism